OpenStudy (anonymous):

expand 3 (× + r + 5)

OpenStudy (jojokiw3):

Use the distributive property of multiplication. a(b + c) = ab + ac.

TheSmartOne (thesmartone):

When you use the distributive property it will become 3(x)+3(r)+3(5). Can you simplify it now? @joeurdaman

TheSmartOne (thesmartone):

Simplify 3(x)=3x 3(r)=3r 3(5)=15 So 3(x)+3(r)+3(5)= 3x+3r+15